OpenStudy (anonymous):

ok last one....I have been trying to solve these couple problems for 3 days...ughh Find the LCM of 25y^2+100y and 5y^2+35y+60

OpenStudy (anonymous):

25y(y + 100) 5(y^2 + 7y + 12) = 5(y + 4)(y+ 3) common factor is 5 lcm = 25(y+100)(y+4)(y+3)